1. A business believes it can make an investment in a piece of machinery that will be used in its manufacturing process for the next 20 years. If the business chooses to purchase the machinery, what financing option would most logically be used to finance the investment without using equity financing?

a) preferred stock

b) 20-year bonds

c) common stock

d) short term note

2. Vladimir Boscak invested $35,000 on March 1, 2009, and in return he received a total of $39,600, which he collected on March 1, 2010. The rate of return on Vladimir's investment was:

a) 7.6%

b) 12.3%

c) 11.6%

d) 13.1%
Answer #1

1. 20- year bonds

Bond financing is more suitable for financing capital assets. Company sells bonds to investors and repay the amount when it is due with interest according to specified schedules.

2. 13.1%

Calculation

Return on investment = 39600-35000= 4600

Rate of return = 4600/ 35000= 13.1%